# Superfluid Dark Matter

### v1

@@ 5.4 Behavior around black holes
Following methodologies outlined in [252, 248], one can determine the DM density profile within this regime by solving the hydrostatic equilibrium equation alongside a heat equation under the gravothermal fluid approximation [169, 253, 252, 254]. These equations are to be solved down to the radius $r_{\text{\tiny mb}}=4GM_{\text{\tiny BH}}$ [255, 252], corresponding to marginally-bound circular orbits within the Schwarzschild geometry. At that point, DM particles are entirely accreted into the BH, resulting in a sharp decline in the DM density profile.
@@ 6 Dynamical friction
For a collisionless medium, the drag force on a linearly moving body has been computed in the pioneering study of Chandrasekhar [259], while the linear motion in a gaseous medium has been discussed in Refs. [260, 261, 168]. The case of circular binaries was investigated in Refs. [262, 263, 264, 265, 266, 267]. More diverse environments, such as fuzzy DM, self-interacting or superrandiance-driven scalar clouds, and Bose-Einstein condensates, have also been explored recently in Refs. [85, 268, 269, 270, 271, 272, 273, 274, 275]. For superfluids—which exhibit frictionless flow—dynamical friction is expected to be significantly suppressed. This feature could lead to intriguing astrophysical implications and potentially resolve several challenges faced by CDM. For instance, the velocity of galactic bars within spiral galaxies is observed to remain nearly constant, contrary to the expected slowdown in CDM [276, 277, 278]. Similarly, the absence of mergers among the five globular clusters orbiting the Fornax dwarf galaxy poses a puzzle: in a cuspy CDM halo, these clusters should have long since migrated to the center due to dynamical friction [279]. Suppressed dynamical friction in a SDM framework could naturally account for their continued survival. See also Ref. [280] for a recent analysis in the context of self-interacting fuzzy DM.
@@ 6.1.3 Explicit evaluation of the friction force
$|F_{\text{\tiny DF}}|=\left\{\begin{array}[]{cl}\frac{4\pi G^{2}M^{2}\rho_{0}}{v^{2}}~{}{\rm log}\Big{(}\frac{mc_{s}R_{\text{\tiny SF}}}{\pi}\sqrt{\mathcal{M}^{2}-1}\Big{)}&\text{for}\hskip 10.0pt\mathcal{M}\geq\sqrt{1+\frac{\pi^{2}}{m^{2}c_{s}^{2}R_{\text{\tiny SF}}^{2}}}\,;\\ \\ 0&\text{for}\hskip 10.0pt\mathcal{M}\leq\sqrt{1+\frac{\pi^{2}}{m^{2}c_{s}^{2}R_{\text{\tiny SF}}^{2}}}\,,\end{array}\right.$ (6.14)
@@ 6.1.4 Accelerated linear motion in the quasi-particle formalism
To provide a simple result, let us consider only the rate of radiating phonons from external lines, induced by the last term in Eq. (6.20), given by $\sim\chi^{2}\Delta^{-1}\chi^{2}$ In principle, one could allow for phonon-mediated scattering or, more generally, for processes where the external phonon line can be attached at more locations besides the external $\chi$-line. However, in the case of soft bremsstrahlung, the Lagrangian is dominated by the operator $\chi^{2}\Delta^{-1}\chi^{2}$, such that the total energy radiated in phonons, as the perturber velocity changes from $v$ to $v^{\prime}$ due to an external field, is [164]
@@
This recovers the same dependence on $G$$M$ and $\rho_{0}$ as the one for a freely-moving perturber. Moreover, contrarily to the logarithmic dependence of the free case, the radiated energy presents an IR divergence $\propto 1/k_{\text{\tiny min}}\sim R_{\text{\tiny SF}}$, and is characterized by a factor $\big{(}\vec{v}-\vec{v}{\,{}^{\prime}}\big{)}^{2}$, as expected for radiation reaction effects, proportional to the square of the acceleration.
@@ 6.2.1 Formalism
This expression gives the perpendicular $(\hat{r})$ and tangential $(\hat{\varphi})$ components of the drag force experienced by a probe moving on a circular orbit. Spherical symmetry dictates that the force along the direction perpendicular to the orbital plane vanishes. The coefficients $\gamma_{\ell m}$ are completely determined by the angular trajectory and given by
@@
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+~{}{\rm i}j_{\ell}\big{(}\ell_{\rm q}\mathcal{M}f^{-}_{m}\big{)}y_{\ell-1}\big{(}\ell_{\rm q}\mathcal{M}f^{-}_{m}\big{)}+\frac{2}{\pi}i_{\ell}\big{(}\ell_{\rm q}\mathcal{M}f^{+}_{m}\big{)}k_{\ell-1}\big{(}\ell_{\rm q}\mathcal{M}f^{+}_{m}\big{)}\biggr{]}\,,$ (6.27)
@@
$f^{\pm}_{m}$ $=\sqrt{2\pm 2\sqrt{1+\left(\frac{m}{\ell_{\rm q}}\right)^{2}}}\quad\Longrightarrow\quad\begin{cases}f_{m}^{+}\simeq 2\,,~{}~{}f^{-}_{m}\simeq\frac{m}{\ell_{\rm q}}\quad~{}~{}~{}~{}~{}\text{for}\quad m\ll\ell_{\rm q}\,;\\ f^{\pm}_{m}\simeq\sqrt{\frac{2m}{\ell_{\rm q}}}\quad~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{for}\quad m\gg\ell_{\rm q}\,,\end{cases}$ (6.28)
@@ 6.2.2 Evolution of black hole binaries and gravitational wave dephasing
To estimate the drag force, we can first appreciate that the light object follows a subsonic motion within the spike, reaching $\mathcal{M}\approx 1$ only when the BH strongly modifies the DM density profile at $r\lesssim 10^{-4}\,r_{h}$ [183]. Furthermore, the quantum-pressure multipole scale $\ell_{\rm q}$ is found to be always much larger than the finite-size cutoff $\ell_{\text{\tiny max}}$ [183]
@@ 7 Emergent phonon-mediated forces in superfluids
While the $\Lambda$CDM model has been remarkably successful at matching cosmological observations on large scales, its explanatory power on galactic scales has been the subject of ongoing debate [68]. Despite the complex role of baryonic feedback processes in galaxy formation, galaxies exhibit a surprising degree of regularity, with striking correlations among their physical properties. Disc galaxies display a remarkably tight correlation between the total baryonic mass (stellar + gas) and the gravitational acceleration in galaxies [305, 306]. This apparent conspiracy, known as the mass discrepancy acceleration relation (MDAR), states that the acceleration experienced by a baryonic particle (irrespective of whether it is due to DM, new long-range forces, or both) can be uniquely predicted from the baryon density profile. At large distances, the MDAR implies the baryonic Tully-Fisher relation [307], which relates the total baryonic mass to the asymptotic/flat rotation velocity as $M_{\rm b}\sim v^{4}_{\rm f}$. This relation holds over five decades in mass, with remarkably small scatter. Another scaling relation is the correlation between the central stellar and dynamical surface densities in disk galaxies [308].
Much work has been done to understand how the MDAR arises within $\Lambda$CDM. Hydrodynamical simulations that generate galaxies whose properties broadly match observations also tend to reproduce the MDAR (e.g.[309]), although with some discrepancy in the trend and scatter [310, 311]. Starting from the early study by [312], various semi-analytical arguments have been put forth in recent years to explain the MDAR in $\Lambda$CDM [313, 314, 315, 316, 317, 318].
@@
$a=\left\{\begin{array}[]{cl}a_{\rm b}\qquad~{}~{}~{}~{}a_{\rm b}\gg a_{0}\\[14.22636pt] \sqrt{a_{\rm b}a_{0}}\qquad a_{\rm b}\ll a_{0}\,,\end{array}\right.$ (7.1)
@@ 7.1 Effective field theory of phonon-mediated force
where $Y=\hat{\mu}_{\text{\tiny NR}}+\dot{\pi}+\vec{v}\cdot\vec{\nabla}\pi$ depends on the velocity $\vec{v}$ of the normal fluid component. The coefficient $\beta$ parametrizes the size of temperature-dependent effects, with $\beta\to 0$ to recover the original action of Eq. (7.6). This parameter should implicitly depend on $T/T_{\rm c}$, even though for simplicity we treat it as a constant. As shown in Refs. [7, 8]$\beta\geq 3/2$ is necessary to reproduce the MDAR with stable perturbations.
@@
$a_{\pi}(r)\simeq\begin{cases}\sqrt{a_{0}a_{\rm b}(r)}\,,\qquad~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{phonon-dominated regime}\,;\\ \sqrt{\frac{3}{2m\hat{\mu}_{\text{\tiny NR}}}}\sqrt{\frac{\beta-1}{2\beta-3}}\alpha^{2}\Lambda a_{\rm b}(r)\,,\quad~{}~{}~{}\text{DM halo-dominated regime}\,,\end{cases}$ (7.12)
@@
A more precise calculation of the accelerations in the two regimes is described in detail in Ref. [220]. There, as an example, the corresponding density profiles were fitted to two representative galaxies, a low surface brightness galaxy IC 2574 and a high surface brightness galaxy UGC 2953, providing an excellent fit to their rotation curves. Phenomenological implications for other astrophysical systems were also discussed in Refs. [220, 357, 358, 164]. [^44] See also Ref. [359] for simulations of structure formation in MONDian SDM models, where strong DM–baryon interactions lead to faster structure growth compared to $\Lambda$CDM, potentially allowing constraints to be placed on the coupling strength to baryons.
@@
Although the idea of a phonon-mediated force may at first seem somewhat exotic, it has a clear physical meaning in the language of fluid mechanics. The idea is captured by the right panel of Fig. 11, which shows the redistribution of DM condensate in the presence of a central baryon lump (large grey disk). There are two possible scenarios, depending on the sign of the coupling between baryons and DM, which determines whether DM is expelled by or accreted onto the baryons. The far away probe baryon (small black disk) can feel the redistribution of the DM condensate. Furthermore, if it couples to DM in the same way as the central lump (e.g., if both are comprised of the same particles) then it will be pulled towards it in both pictures.
Figure 11: Left panel: Galactic forces between baryons. The light blue region depicts the envelope of the normal DM component, enclosing the superfluid core (blue area), while brown discs depict baryons. Wavy lines correspond to the gravitational interaction sourced both by baryons and DM, while the dashed line depicts the phonon-mediated force. Right panel: Physical picture of environmentally emergent forces between baryons (gray and black disks) as a result of the displacement of the DM condensate (light blue).
@@ 7.3 Emergent long-range interaction from $U(1)$-invariant coupling
[…]
@@
Evidently, the force is an inverse-square law $\sim 1/r^{2}$ for $r\ll\ell$, and becomes long-range ($\ell\rightarrow\infty$) in the limit of vanishing self-interaction, $\lambda_{4}\rightarrow 0$.
@@ 7.3.1 Distorted Bose-Einstein condensates
[…]
@@
[…]
@@
$F_{h}=\begin{cases}\frac{v^{2}}{4\pi\Lambda^{4}}\frac{M_{1}M_{2}}{r^{2}}\,,\qquad\text{unscreened source and probe}\,;\\ 4\pi v^{2}\frac{R_{1}R_{2}}{r^{2}}\,,\qquad\,\text{screened source and probe}\,;\\ \frac{v^{2}}{\Lambda^{2}}\frac{R_{1}M_{2}}{r^{2}}\,,\qquad\,\,\,\,\,\,\text{screened source 1 and unscreened probe 2}\,.\end{cases}$
@@ 7.3.2 Bounds on emergent forces: low-surface brightness galaxies
[…]
@@
[…]
@@ 8 Conclusions
Building on this foundation, Secs. 3 and 4 addressed the cosmological evolution of the model and the conditions under which superfluid soliton states can emerge in galaxies. Dark matter production proceeds non-thermally via the vacuum misalignment mechanism, leading to a highly homogeneous and coherent scalar field profile in the early universe. As the Hubble parameter redshifts to the scale of the DM mass, the field begins to behave like a condensate of massive particles, eventually forming halos through virialization. In the early stages of virialization, when self-interactions are negligible, the system gradually loses its initial coherence and behaves as cold, pressureless matter. However, as self-interactions become important, they enable re-equilibration and the restoration of coherence, culminating in the formation of a superfluid phase within galaxies. This state evolves through gravitational interactions, eventually yielding a dense central core surrounded by streams of superfluid debris.
@@
The formation of superfluid cores in galaxies has profound implications for their structure and dynamics, as discussed in Secs. 5--7. These include modified cluster merger dynamics—leading to a revised interpretation of the Bullet Cluster bound—and the formation of vortex lattices that preserve angular momentum through the phase transition. Additionally, superfluid dark matter generates density spikes around supermassive black holes in the center of galaxies, offering distinctive observational signatures that differentiate this model from other DM candidates. We also examined the behavior of binary systems embedded in a superfluid background, deriving predictions for dynamical friction and assessing prospects for detection in future gravitational wave experiments. Finally, the interaction of phonons with external sources, such as baryons, gives rise to long-range forces which, in some scenarios, reproduce MOND-like phenomenology and the associated galactic scaling relations.
In summary, the study of superfluid dark matter is a rich and rapidly evolving research program, with numerous theoretical, numerical, and observational opportunities for further development. Foremost among these is the necessity of conducting sophisticated numerical simulations to fully understand the formation and evolution of superfluid structures across cosmic time. Such simulations are essential to accurately compare the model to data, and may reveal novel observational signatures of this model on both small and large scales. Equally important is constructing a consistent particle physics framework for these ultralight bosons, encompassing their cosmological origin and thermal history. Lastly, a systematic evaluation of the model’s ability to address the small-scale challenges to the $\Lambda$CDM model is crucial for assessing its viability. We hope this review provides a useful overview of the current state of the field and offers a springboard for future investigations.

### v2

@@ 5.6 Behavior around black holes
Following methodologies outlined in [409, 141], one can determine the DM density profile within this regime by solving the hydrostatic equilibrium equation alongside a heat equation under the gravothermal fluid approximation [35, 319, 409, 282]
$\frac{{\rm d}T}{{\rm d}r}=\frac{D}{\kappa r^{2}(1-2GM_{\text{\tiny BH}}/r)^{3/2}}-\frac{T}{1-2GM_{\text{\tiny BH}}/r}\frac{GM_{\text{\tiny BH}}}{r^{2}}\,$, (5.35)
which has been previously introduced in Eq. (3.12) of Sec. 3.2 for the case of self-interacting DM. This equation relates the temperature gradients of a virialized gas with temperature $T=\frac{1}{3}mv^{2}$ and thermal conductivity $\kappa$, at rest in a stationary and spherically-symmetric gravitational field, with the total radiated heat flux, assumed to be constant (and proportional to the constant $D$) for a steady-state solution [34]. The expression for the thermal conductivity $\kappa$ is obtained by matching the relativistic heat flux equation to the Newtonian result of kinetic theory [409], giving
$\kappa\simeq\frac{\sqrt{3}}{2}AC\frac{\sigma\rho^{2}r^{2}v}{m^{2}}\,$, (5.36)
in the long mean free path regime $\frac{m}{\rho\sigma}\gg r$ [141], where $\sigma$ is the scattering cross section, $C\simeq\frac{58}{77}$ is a constant determined by N-body simulations [282], and $A=\sqrt{\frac{16}{\pi}}$ for hard-sphere interactions.
These equations are to be solved down to the radius $r_{\text{\tiny mb}}=4GM_{\text{\tiny BH}}$ [397, 409], corresponding to marginally-bound circular orbits within the Schwarzschild geometry. At that point, DM particles are entirely accreted into the BH, completely determining the heat constant $D$ and resulting in a sharp decline in the DM density profile.
@@ 6 Dynamical friction
For a collisionless medium, the drag force on a linearly moving body has been computed in the pioneering study of Chandrasekhar [123], while the linear motion in a gaseous medium has been discussed in Refs. [396, 387, 365]. The case of circular binaries was investigated in Refs. [276, 277, 296, 297, 151, 101]. More diverse environments, such as fuzzy DM, self-interacting or superradiance-driven scalar clouds, and Bose-Einstein condensates, have also been explored recently in Refs. [250, 289, 20, 226, 452, 86, 87, 47, 432]. For superfluids—which exhibit frictionless flow—dynamical friction is expected to be significantly suppressed. This feature could lead to intriguing astrophysical implications and potentially resolve several challenges faced by CDM. For instance, the velocity of galactic bars within spiral galaxies is observed to remain nearly constant, contrary to the expected slowdown in CDM [144, 145, 395]. Similarly, the absence of mergers among the five globular clusters orbiting the Fornax dwarf galaxy poses a puzzle: in a cuspy CDM halo, these clusters should have long since migrated to the center due to dynamical friction [435]. Suppressed dynamical friction in a SDM framework could naturally account for their continued survival. See also Ref. [298] for a recent analysis in the context of self-interacting fuzzy DM.
@@ 6.1.3 Explicit evaluation of the friction force
[…]
@@ 6.1.4 Accelerated linear motion in the quasi-particle formalism
To provide a simple result, let us consider only the rate of radiating phonons from external lines, induced by the last term in Eq. (6.20), given by $\sim\chi^{2}\Delta^{-1}\chi^{2}$ . In principle, one could allow for phonon-mediated scattering or, more generally, for processes where the external phonon line can be attached at more locations besides the external $\chi$-line. However, in the case of soft bremsstrahlung, describing the emission of low-energy (soft) particles from an accelerated or scattered perturber, the Lagrangian is dominated by the operator $\chi^{2}\Delta^{-1}\chi^{2}$, such that the total energy radiated in phonons, as the perturber’s velocity changes from $\vec{v}$ to $\vec{v}{\,{}^{\prime}}$ due to an external field, is [60]
@@
This recovers the same dependence on GGMM and $\rho_{0}$ as the one for a freely-moving perturber. Moreover, contrarily to the logarithmic dependence of the free case, the radiated energy presents an IR divergence $\propto 1/k_{\text{\tiny min}}\sim R_{\text{\tiny SF}}$, and is characterized by a factor $\big(\vec{v}-\vec{v}{\,{}^{\prime}}\big)^{2}$, as expected for radiation reaction effects, proportional to the square of the acceleration.
@@ 6.2.1 Formalism
[…]
@@
[…]
@@
[…]
@@ 6.2.2 Evolution of black hole binaries and gravitational wave dephasing
To estimate the drag force, we can first appreciate that the light object follows a subsonic motion within the spike, reaching $\mathcal{M}\approx 1$ only when the BH strongly modifies the DM density profile at $r\lesssim 10^{-4}\,r_{h}$ [55], where $r_{h}$ has been defined as the BH sphere of influence in Eq. (5.28). Furthermore, the quantum-pressure multipole scale $\ell_{\rm q}$ is found to be always much larger than the finite-size cutoff $\ell_{\text{\tiny max}}$ [55]
@@ 7 Emergent phonon-mediated forces in superfluids
[…]
@@
[…]
@@ 7.1 Effective field theory of phonon-mediated force
[…]
@@
where primes denote radial derivatives, and $a_{\rm b}(r)=\frac{GM_{\rm b}(r)}{r^{2}}$. This implies a cubic equation for $\pi^{\prime\,2}$, whose real root does not have a particularly illuminating analytic form. Instead, it is useful to consider different limiting regimes, following [63, 64]. Sufficiently close to the baryon source (i.e., in the central regions of galaxies), one has $\pi^{\prime\,2}\gg 2m\hat{\mu}_{\text{\tiny NR}}$, such that the desired behavior is obtained, $\pi^{\prime}=\sqrt{\alpha M_{\rm Pl}a_{\rm b}(r)}$. In the opposite regime far from baryons, $\pi^{\prime\,2}\ll 2m\hat{\mu}_{\text{\tiny NR}}$, the solution instead scales like $\pi^{\prime}=\sqrt{\frac{3}{2m\hat{\mu}_{\text{\tiny NR}}}}\sqrt{\frac{\beta-1}{2\beta-3}}\alpha M_{\rm Pl}a_{\rm b}(r)$ for approximately constant $\hat{\mu}_{\text{\tiny NR}}$. Thus the $\pi$-mediated acceleration takes the form
@@
[…]
@@
Although the idea of a phonon-mediated force may at first seem somewhat exotic, it has a clear physical meaning in the language of fluid mechanics. The idea is captured by the right panel of Fig. 11, which shows the redistribution of DM condensate in the presence of a central baryon lump (large gray disk). There are two possible scenarios, depending on the sign of the coupling between baryons and DM, which determines whether DM is expelled by or accreted onto the baryons. The far away probe baryon (small black disk) can feel the redistribution of the DM condensate. Furthermore, if it couples to DM in the same way as the central lump (e.g., if both are comprised of the same particles) then it will be pulled towards it in both pictures.
Figure 11: Left panel: Galactic forces between baryons. The light blue region depicts the envelope of the normal DM component, enclosing the superfluid core (blue area), while brown disks depict baryons. Wavy lines correspond to the gravitational interaction sourced both by baryons and DM, while the dashed line depicts the phonon-mediated force. Right panel: Physical picture of environmentally emergent forces between baryons (gray and black disks) as a result of the displacement of the DM condensate (light blue).
@@ 7.3 Emergent long-range interaction from $U(1)$-invariant coupling
To derive the force law, let us focus on static sources, setting $\dot{h}=\dot{\pi}=0$. [^51] For a real scalar field undergoing Bose-Einstein condensation, the long-range force would be characterized by an additional rapid oscillatory prefactor [142] that can be removed by averaging over time. The Lagrangian reduces to
@@
Evidently, the force is an inverse-square law $\sim 1/r^{2}$ for $r\ll\ell$, it is exponentially suppressed at distances greater than the healing length, where self-interactions start to smooth out the perturbation induced in the condensate by the localized source, and it becomes long-range ($\ell\rightarrow\infty$) in the limit of vanishing self-interaction, $\lambda_{4}\rightarrow 0$.
@@ 7.3.1 Distorted Bose-Einstein condensates
$\left.M_{\rm eff}\right|_{x\ll 1}\simeq\frac{4\pi}{3}\rho R^{3}\equiv M\qquad\longrightarrow\qquad h_{\rm out}(r)=-\frac{v}{\Lambda^{2}}\frac{M}{4\pi r}\qquad(\text{low-density/unscreened})\,.$ (7.28)
@@
$\left.M_{\rm eff}\right|_{x\gg 1}\simeq 4\pi\Lambda^{2}R\qquad\longrightarrow\qquad h_{\rm out}(r)=-v\frac{R}{r}\qquad(\text{high-density/screened})\,$, (7.29)
@@
[…]
@@ 7.3.2 Bounds on emergent forces: low-surface brightness galaxies
$m\lesssim 3\times 10^{-23}~{\rm eV}\cdot\left(\frac{\Lambda}{10^{-4}M_{\rm Pl}}\right)^{-2}\,;\qquad\text{linear/unscreened regime}\,$, (7.31)
@@
$m\lesssim 2\times 10^{-23}~{\rm eV}\cdot\left(\frac{\Lambda}{10^{-4}M_{\rm Pl}}\right)^{-1}\,;\qquad\text{non-linear/screened regime}\,.$ (7.32)
@@ 8 Discussions and conclusions
Building on this foundation, Secs. 3 and 4 addressed the cosmological evolution of the model and the conditions under which superfluid soliton states can emerge in galaxies. Despite the lack of a concrete UV completion of the proposed scalar field model, it is plausible to assume that sub-eV DM particles could be produced non-thermally via the vacuum misalignment mechanism, which would generically lead to a highly homogeneous and coherent scalar field profile in the early universe. As the Hubble parameter redshifts to the scale of the DM mass, the field begins to behave like a condensate of massive particles, eventually forming halos through virialization. In the early stages of virialization, when self-interactions are negligible, the system gradually loses its initial coherence and behaves as cold, pressureless matter. However, as self-interactions become important, they enable re-equilibration and the restoration of coherence, culminating in the formation of a superfluid phase within galaxies. This state evolves through gravitational interactions, eventually yielding a dense central core surrounded by streams of superfluid debris.
@@
The formation of superfluid cores in galaxies has profound implications for their structure and dynamics, as discussed in Secs. 5-7. These include modified cluster merger dynamics—leading to a revised interpretation of the Bullet Cluster bound—and the formation of vortex lattices that preserve angular momentum through the phase transition. Additionally, while various dark matter models produce density spikes around supermassive black holes in galactic centers, superfluid dark matter generates a characteristic spike with distinctive observational features that set it apart from other DM candidates. We also examined the behavior of binary systems embedded in a superfluid background, deriving predictions for dynamical friction and assessing prospects for detection in future gravitational wave experiments. Finally, the interaction of phonons with external sources, such as baryons, gives rise to long-range forces which, in some scenarios, reproduce MOND-like phenomenology and the associated galactic scaling relations.
In a nutshell, the superfluid dark matter scenario sits at the crossroads of other paradigms, such as cold, warm, self-interacting, and fuzzy dark matter models. It attempts to merge several of their key features under a single framework. On cosmological scales, it reproduces the successful predictions of CDM, behaving effectively as a collisionless fluid and preserving the observed large-scale structure. However, on smaller, galactic scales, the coherence is lost outside the central superfluid soliton due to the combination of the Jeans instability and the tidal disruption of would-be superfluid solitons. Within the central soliton, the superfluid phase transition gives rise to collective excitations that could mediate additional long-range forces between baryons, leading to cored halo profiles and scaling relations that could match galactic observations. Because of the finite Jeans scale, this scenario erases substructures below that characteristic length, in analogy with warm and fuzzy dark matter, which suppress structure below the de Broglie wavelength inside galaxies.
At the same time, this suppression of small-scale power introduces both predictive advantages and potential challenges. The erasure of substructures below the Jeans scale can help mitigate long-standing small-scale issues of CDM, such as the overabundance of satellites and overly dense halo centers. Yet, if the Jeans length is too large, it risks conflicting with the observed abundance and internal dynamics of dwarf galaxies, requiring fine-tuning of the boson mass and self-interaction strength accordingly. However, the physical origin and extent of this suppression differ substantially from other models. In fuzzy dark matter, and to some extent in warm dark matter, the de Broglie wavelength of the particles sets a universal cutoff, preventing the formation of substructures below that scale throughout the halo. In contrast, for superfluid dark matter, this suppression applies only within the thermalized superfluid core, where condensation and collective behavior occur. The outer regions of halos, shaped by tidal disruption of would-be superfluid solitons, consist of virialized superfluid streams that never fully re-thermalize due to the absence of a dense, gaseous dark matter phase with a highly occupied phase space. As a result, these non-thermalized envelopes may be able to form smaller substructures than the Jeans length. However, this qualitative discussion needs to be substantiated by numerical simulations. This dual structure, superfluid core plus virialized outer halo, places superfluid dark matter in a unique middle ground between collisionless and wave-like models.
[…]